OpenStudy (anonymous):

a yard produces two special wines, a white and a red. a bottle of the white wine requires 14 pounds of grapes and 1 hour of processing time. a bottle of red wine requires 25 pounds of grapes and 2 hours of processing time. the vineyard has on hand 2,198 pounds of grapes and can allot 160 hours of processing time to the production of these wines. a bottle of the white wine sells for $11.00 while a bottle of the red wine sells for $20.00. how many bottles of each type should the vineyard produce in order to maximize gross sales?

OpenStudy (anonymous):

help

OpenStudy (anonymous):

Let w represent white wine and r represent red wine we need to equations to solve for these two variable, one will be an equation of total hours and the other of total pounds of grapes Total hours=160 so 1*w+2*r=160 Total weight=2198 so 14*w+25*r=2198 now solve for one variable in one equation and substitute that expression into the other equation to solve for both variables

OpenStudy (anonymous):

still lost

OpenStudy (anonymous):

so w=160-2r now plug that expression for w into the other equation and you can solve for the number of bottle of red wine

OpenStudy (anonymous):

if you let me know exactly what your stuck on i can try to help with it

OpenStudy (anonymous):

what i plug into i understand the equations but dont know what i plug in

OpenStudy (anonymous):

so w=160-2r and 14*w+25*r=2198 which is equal to 14*(160-2r)+25*r=2198 because we solved the other equation in terms of w we can plug that into the second equation now that equation is reduced to only one variable which we can solve for

OpenStudy (anonymous):

once we know the value for r we can plug it into one of the equations and solve for w